# Modeling portfolio loss distribution under infectious defaults and immunization

### v3

@@ 4.2 Mixture contagion loss distribution
In this work, for the “correlated default state” we opted for a simple one factor Gaussian model as it is easy to calibrate, and well known among regulators and practitioners, but we could choose an alternative model that accounts for fatter tails such as the one proposed in Kalemanova et al., (2007) that assumes a NIG distribution if the fit of the model is not satisfactory. As discussed in Section 5.2, we consider a restricted version of the model with three parameters that allows us to satisfactory calibrate the quotes of iTraxx Synthetic CDO index tranches.
We underline that mixture models are commonly used in the modelization of portfolio credit risk (see e.g. CreditRisk+, Suisse,, 1997), although in a different way: Frey and McNeil, (2003) extensively discuss Bernoulli mixture models in which defaults present a conditional independence structure conditional on common factors, showing how latent variable models such as KMV (Kealhofer and Bohn,, 2001) and RiskMetrics (Morgan et al.,, 1997) can be mapped to equivalent mixture models. In such cases, the conditional distribution of portfolio losses is integrated on the value of the conditioning factors. In our approach, instead, the loss distribution in each of the two possible states of the world (“contagion” and “correlated default”) is integrated on the distribution of the mixing binary variable $\xi$.
Conceptually, our two-state approach is related to the literature on regime-switching models, that are statistical frameworks in which parameters or dynamics change according to an unobserved state variable that evolves over time, typically governed by a Markov process. such models have been used in systemic risk analysis (see e.g. Abdymomunov,, 2013; Bernardi and Catania,, 2019) to address some observed market characteristics. In particular, the literature documents the presence of structural breaks in the risk transmission in crises periods (Billio et al.,, 2012) and default clustering (Berndt et al.,, 2010; Azizpour et al.,, 2018; Montagna et al.,, 2020). Regime switching models have also been widely used in credit risk measurement (see e.g. Davies,, 2004; Liang and Wang,, 2012). An approach similar to ours is the one proposed by Davidson, (2020), who develops a Bayesian time-varying parameter VAR model to distinguish between interdependence, contagion via interdependence, and abrupt contagion using a model-switching framework in the context of Latin American crises. Although her focus is on macro‑financial linkages, this typology shares the conceptual intuition behind our two‑regime structure. In contrast, our model explicitly differentiates default dynamics through a single-factor model during peace periods and a contagion‑driven model during crises.
@@ 5.2 Specifications and calibration of the loss distribution
The OFG model is a single-parameter model widely used in the industry for its simplicity and tractability. It is also considered a lingua franca among practitioners: the OFG base correlation is used on CDO trading desks in much the same way the Black-Scholes implied volatility is used by option traders. Readers unfamiliar with the OFG model might refer to Li, (2000) for the original description or to Burtschell et al., (2009) for a good review of the methodology. Let just remember here that in the OFG framework a single correlation input $\rho\in[0,1]$ describes the likelihood of names to default together. In practice, though, it is well known that the model with a single parameter $\rho$ cannot explain the prices of the many tranches one observes on the market. [^1] The standard solution is to resort to the base correlation approach. Every tranche is priced as a difference of two equity tranches, i.e. tranches with attachment set at zero (for example, a $3\%-6\%$ tranche will be valued as the difference of a $0\%-6\%$ tranche minus a $0\%-3\%$ one). These two valuations are treated as independent from each other and hence different specification of the model are used (i.e. different values of $\rho$ are employed). This approach, known as base correlation, has the drawback of creating the potential for mispricing for non-standard tranches. For simplicity we assume uniform default probabilities, that are set equal to the average of the default probabilities $\tilde{p}_{i}(t)$ estimated from the single name CDS spreads of the index contituents.
@@
  • The quantities $v_{i}(t)$ are in a monotonic decreasing relationship with $\tilde{p}_{i}(t)$; this reflects the fact that healthier firms have a bigger impact in case of idiosyncratic default than riskier ones; the market is expecting default of risky firms (hence the high probability of default) and therefore the shock when the event finally happen is minor. We restrict the parameter $\mu_{i}$ to be either constant across all the names, or differentiated according to industrial sector, obtaining three variants of the model (see below);
@@ 5.2.1 Choice of the parameters $\mu_{i}$
As discussed above, we limit the calibration of the contagion model to one parameter, while maintaining a certain degree of heterogeneity and using the information relative to the idiosyncratic default probabilities $\tilde{p}_{i}(t)$. We can further tune the model by restricting the values of $\mu_{i}$ depending on economic considerations. We consider three alternatives: first we set the value constant for all the names in the index (model CON-FLAT). In particular, we set the value to $\mu_{i}=0.1$ for all $i$. We then consider two alternative calibrations that attribute a larger potential for the transmission of contagion to banking and financial institutions, as the literature on systemic risk highlights the centrality of such sectors in the amplification of risk and the transmission of financial distress to the whole economic system (Schwarcz,, 2008; Freixas et al.,, 2015; Montagna et al.,, 2020). In particular, we consider the model CON-BNK that set the coefficient $\mu_{i}=0.2$ for companies in the banking sector and $\mu_{i}=0.05$ for other companies, and the model CON-FIN that set the coefficient $\mu_{i}=0.2$ for companies in the banking and financial sector and $\mu_{i}=0.05$ for others. The same applies to the mix and conditional models, for which we consider the three variants each: (COND-FLAT), (COND-BNK), (COND-FIN), (MIX-FLAT), (MIX-BNK), and (MIX-FIN). Alternatively, we could have also parametrized $\mu_{i}$ to some market based measure of systemic relevance, such as the the connectedness measures computed by network measures computed by Diebold and Yilmaz, (2012) using volatility spillover networks. Such measures however are typically computed on historical data, while we wanted to use forward looking measures.
In preliminary analyses we also tested richer parametrizations for the contagion models that included one or two additional parameters to control the size of $\mu_{i}$ across sectors, but we maintained the proposed parametrization due to the satisfactory empirical fit, and model parsimony. In Appendix C we test the pricing of iTraxx tranches using alternative values of $\mu_{i}$ as a robustness check.
@@ 6 Simulation study
Figure 1: Sensitivity of the loss distribution to parameters $\tilde{p}$, $\omega$, $\mu$. The baseline parameters for the distribution are $\tilde{p}=0.05$, $\omega=0.5$, $\mu=0.1$, and there are 125 names in the portfolio. The horizontal axis is capped at 40 to make the visualization clearer. Parameters are homogeneous across all names.
We see that overall, the resulting distributions are bimodal, in line with the presence of two distinct possible outcomes depending on the presence or absence of an active infection. Specifically, in the first panel we see the effect of the parameter $\tilde{p}$, finding that the increase of the marginal default probability shifts the distribution to the right. In the middle panel we notice that an increase of the value of $\omega$ amplifies the bimodality of the distribution separating the peaks. Finally, on the right we see that changes in the value of $\mu$ affect mostly the second peak of the distribution without affecting significantly the main peak. Overall, the contagion model allows to account for flexible shapes of the loss distribution. We point out that there is the potential for identifiability issues do to the similar effects of the parameters $\mu$ and $\omega$ on the loss curve. Thus, in the empirical analysis, $\mu_{i}$ was held constant to ensure model parsimony.
@@
Next, we compare the loss distributions under different models: pure contagion (CON), one factor Gaussian copula (OFG), conditional model (COND), and mixture model (MIX). To make the comparison fair, we set the parameters of each distribution in such a way that the expected losses (EL) and the pairwise default correlation between any two assets ($\rho(Z_{i},Z_{j})$) are the same for each model. LGD is set to 100%, and the parameters for each distribution are reported in Table 1. In particular, we set a baseline parametrization for the CON model, and we adjusted the correlation parameter $\rho$ in the models. Default correlation is computed for CON, COND, and MIX models as described in Section 4.3, while for the OFG model we computed them analytically as described e.g. in Vasicek, (2002).
@@
Figure 2: Loss distribution for different models with the parameters in Table 1. The horizontal axis is capped at 40 to make the visualization clearer.
@@ 7.1 Dataset
We consider synthetic CDO tranches of the iTraxx Europe index, that is the main reference for the European market and it is the largest in terms of CDS index tranches [^2] The notional traded in iTraxx Europe index tranches in 2024 was USD 99.1 billion, ranking above the CDX North America Investment Grade, despite the decrease in volumes compared to the previous year (see S&P Dow Jones Indices - 2024 & H1 2025 Fixed Income Index Products Report, https://www.spglobal.com/spdji/en/documents/commentary/market-commentary-fixed-income-index-products-annual-report-2024-2025-h1.pdf). Tranches are quoted in terms of upfront paid by the protection buyer to enter in a contract with 100 basis point coupon. The main index is instead listed in terms of par spread (i.e. the level of the premium that balances the expected value of the two legs of the contract). The quotes for contracts with 5 years maturity have been downloaded from LSEG Workspace for the index and for four tranches: Equity (0-3%), Junior Mezzanine (3-6%), Senior Mezzanine (6-12%), Senior (12-100%). For each day we also download from LSEG Workspace the local 5Y CDS spread for the 125 constituents of the index to compute the marginal default probabilities (used to estimate the marginal default probabilities $\tilde{p}_{i}$), the industry sector to define $\mu_{i}$ (for CON-BNK we consider companies whose industry sector is “Banking”, while for CON-FIN we consider all the companies belonging to the “Banking”, “Finance”, or “Insurance” industry sectors), and the bootstrapped EURIRS swap rates as the risk free spot rates. We assume a 40% recovery rate. Every 6 months a new series is released, so that the index composition reflects the market characteristics. We download the dataset with daily frequency from October 31, 2019 to May 30, 2025, and for each date we consider the most recent iTraxx series available. Overall we considered 1457 trading days, excluding 2.8% of the dates from the analysis due to missing quotes for the CDO tranches.
@@ 7.2 Pricing with contagion and mix models
For the calibration of the model we use MATLAB 2024a, minimizing an objective function consisting in the sum of root squared percentage error of the quotes (upfronts for the tranches, par-spread for the whole index). Since some of the quotes were very close to zero, we added a translation of 0.1 in the quotes used in the objective function to avoid denominator close or equal to zero to improve the numerical stability of the optimization. The objective function is minimized using fmincon from the Optimization Toolbox, and the initial parameters supplied to the solver are $\omega_{0}=\rho_{0}=\pi_{0}=0.5$. Each optimization of the COND model (the most computationally intensive, since the contagion loss distribution must be computed for each value of the conditioning variable) takes typically less than two minutes (single core) on a Macbook air with M1 processor and 16 GB of RAM. We bound all the parameters between 0.05 and 0.95 to improve the numerical stability, and, with the exception of the OFG model, in the majority of times we do not obtain corner solutions. [^3] The code is available from the corresponding author upon request.
@@
Table 3: iTraxx tranche calibrated quotes. The quotes for the CDO tranches are upfronts (running spread equal to 100 bps) while the ones for the index are par spreads.
@@
To study the models’ adaptive boundaries in different market environments, we report the average error in a set of crisis scenarios. In particular, we consider the following 6 scenarios: COVID-19 Sharp Shock (Mar 1 – Jun 1, 2020), COVID-19 Recovery & Adaptation (Jun 1 – Dec 31, 2020), Post-Pandemic (Jan 1 – Dec 31, 2021), Ukraine / Energy Crisis (Feb 24 – Aug 31, 2022), European Interest Rate Tightening (Jul 27 – Oct 27, 2022), Post-Crisis Normalization (Jan 1, 2023 – May 31, 2025), and we fit the model each day. Table 4 reports the average MAE of the days in each period and the corresponding volatility. We see that the MIX models are the best performing in each scenario. As we previously noticed from Figure 3, the OFG tends to perform relatively better in quiet periods, while the contagion models (CON) show lower errors in the harshest periods. Although it is outside the scope of the work, we point out that such observation may have implications on the interpretation of the model: the market perception of the loss curve may be closer to the CON model in distressed period with high perceived probability of default clustering, while it may be closer to the one implied by the Gaussian model in quiter periods. Thus, the proposed parametrization of the models may be used to gauge the market perception of the stability of the market: a higher-than-usual level of $\pi$ in the MIX model, may for instance be studied as a market-based indicator of an increased fear of financial contagion.
@@ 8 Conclusions
In this paper we presented a new contagion model for the estimation of the loss distribution of a credit portfolio. The model that we introduce is related to the seminal work of Davis and Lo, (2001) and it can achieve reasonable performance with heterogeneous portfolios.
We then introduced two alternative models for the construction of the loss distribution that incorporate the effect of a common factor, either using a mixture distribution or by conditioning on a common Gaussian factor. We apply the model to the problem of pricing synthetic CDO tranches of the iTraxx index, calibrating several specifications of the model on multiple dates with satisfactory performance. In particular the best performance are obtained using the three parameters mixture models. The models are computationally efficient, and the optimal parameters have a clear economic interpretation.
In addition to fixed-income trading desks, the model has potential applications in the monitoring of systemic risk, as it allows to extract market-based forward looking information on the possible manifestation of systemic events, distinguishing between contagion events and joint defaults due to common factors.

### v4

@@ 4.2 Mixture contagion loss distribution
In this work, for the “correlated default state” we opted for a simple one factor Gaussian model as it is easy to calibrate, and well known among regulators and practitioners, but we could choose an alternative model that accounts for fatter tails such as the one proposed in Kalemanova et al. (2007) that assumes a NIG distribution if the fit of the model is not satisfactory. As discussed in Section 5.2, we consider a restricted version of the model with three parameters that allows us to satisfactory calibrate the quotes of iTraxx Synthetic CDO index tranches.
We underline that mixture models are commonly used in the modelization of portfolio credit risk (see e.g. CreditRisk+, Suisse, 1997), although in a different way: Frey and McNeil (2003) extensively discuss Bernoulli mixture models in which defaults present a conditional independence structure conditional on common factors, showing how latent variable models such as KMV (Kealhofer and Bohn, 2001) and RiskMetrics (Morgan and others, 1997) can be mapped to equivalent mixture models. In such cases, the conditional distribution of portfolio losses is integrated on the value of the conditioning factors. In our approach, instead, the loss distribution in each of the two possible states of the world (“contagion” and “correlated default”) is integrated on the distribution of the mixing binary variable $\xi$.
Conceptually, our two-state approach is related to the literature on regime-switching models, that are statistical frameworks in which parameters or dynamics change according to an unobserved state variable that evolves over time, typically governed by a Markov process. such models have been used in systemic risk analysis (see e.g. Abdymomunov, 2013; Bernardi and Catania, 2019) to address some observed market characteristics. In particular, the literature documents the presence of structural breaks in the risk transmission in crises periods (Billio et al., 2012) and default clustering (Berndt et al., 2010; Azizpour et al., 2018; Montagna et al., 2020). Regime switching models have also been widely used in credit risk measurement (see e.g. Davies, 2004; Liang and Wang, 2012). An approach similar to ours is the one proposed by Davidson (2020), who develops a Bayesian time-varying parameter VAR model to distinguish between interdependence, contagion via interdependence, and abrupt contagion using a model-switching framework in the context of Latin American crises. Although her focus is on macro‑financial linkages, this typology shares the conceptual intuition behind our two‑regime structure. In contrast, our model explicitly differentiates default dynamics through a single-factor model during peace periods and a contagion‑driven model during crises.
@@ 5.2 Specifications and calibration of the loss distribution
The OFG model is a single-parameter model widely used in the industry for its simplicity and tractability. It is also considered a lingua franca among practitioners: the OFG base correlation is used on CDO trading desks in much the same way the Black-Scholes implied volatility is used by option traders. Readers unfamiliar with the OFG model might refer to Li (2000) for the original description or to Burtschell et al. (2009) for a good review of the methodology. Let just remember here that in the OFG framework a single correlation input $\rho\in[0,1]$ describes the likelihood of names to default together. In practice, though, it is well known that the model with a single parameter $\rho$ cannot explain the prices of the many tranches one observes on the market. [^1] The standard solution is to resort to the base correlation approach. Every tranche is priced as a difference of two equity tranches, i.e. tranches with attachment set at zero (for example, a $3\%-6\%$ tranche will be valued as the difference of a $0\%-6\%$ tranche minus a $0\%-3\%$ one). These two valuations are treated as independent from each other and hence different specification of the model are used (i.e. different values of $\rho$ are employed). This approach, known as base correlation, has the drawback of creating the potential for mispricing for non-standard tranches. For simplicity we assume uniform default probabilities, that are set equal to the average of the default probabilities $\tilde{p}_{i}(t)$ estimated from the single name CDS spreads of the index contituents.
@@
  • The quantities $v_{i}(t)$ are in a monotonic decreasing relationship with $\tilde{p}_{i}(t)$; this reflects the fact that healthier firms have a bigger impact in case of idiosyncratic default than riskier ones; the market is expecting default of risky firms (hence the high probability of default) and therefore the shock when the event finally happen is minor. We restrict the parameter $\mu_{i}$ to be either constant across all the names, or differentiated according to industrial sector, obtaining three variants of the model (see below); the square-root transformation applied to $\tilde{p}_{i}(t)$ amplifies variation in the lower range of values and redistributes mass across the support. This choice performs well empirically and alternative concave transformations yield similar results;
@@ 5.2.1 Choice of the parameters $\mu_{i}$
As discussed above, we limit the calibration of the contagion model to one parameter, while maintaining a certain degree of heterogeneity and using the information relative to the idiosyncratic default probabilities $\tilde{p}_{i}(t)$. We can further tune the model by restricting the values of $\mu_{i}$ depending on economic considerations. We consider three alternatives: first we set the value constant for all the names in the index (model CON-FLAT). In particular, we set the value to $\mu_{i}=0.1$ for all $i$. We then consider two alternative calibrations that attribute a larger potential for the transmission of contagion to banking and financial institutions, as the literature on systemic risk highlights the centrality of such sectors in the amplification of risk and the transmission of financial distress to the whole economic system (Schwarcz, 2008; Freixas et al., 2015; Montagna et al., 2020). In particular, we consider the model CON-BNK that set the coefficient $\mu_{i}=0.2$ for companies in the banking sector and $\mu_{i}=0.05$ for other companies, and the model CON-FIN that set the coefficient $\mu_{i}=0.2$ for companies in the banking and financial sector and $\mu_{i}=0.05$ for others. The same applies to the mix and conditional models, for which we consider the three variants each: (COND-FLAT), (COND-BNK), (COND-FIN), (MIX-FLAT), (MIX-BNK), and (MIX-FIN). Alternatively, we could have also parametrized $\mu_{i}$ to some market based measure of systemic relevance, such as the the connectedness measures computed by network measures computed by Diebold and Yilmaz (2012) using volatility spillover networks. Such measures however are typically computed on historical data, while we wanted to use forward looking measures.
In preliminary analyses we also tested richer parametrizations for the contagion models that included one or two additional parameters to control the size of $\mu_{i}$ across sectors, but we maintained the proposed parametrization due to the satisfactory empirical fit, and model parsimony. In Appendix C we test the pricing of iTraxx tranches using alternative values of $\mu_{i}$ to assess the effect of different parameter choices on pricing errors.
@@ 6 Simulation study
Figure 1: Sensitivity of the loss distribution to parameters $\tilde{p}$, $\omega$, $\mu$. The baseline parameters for the distribution are $\tilde{p}=0.05$, $\omega=0.5$, $\mu=0.1$, and there are 125 names in the portfolio. The horizontal axis represents the number of defaults and, although the maximum possible number is 125, only the range up to 40 is displayed for clarity. Parameters are homogeneous across all names.
We see that overall, the resulting distributions are bimodal, in line with the presence of two distinct possible outcomes depending on the presence or absence of an active infection. Specifically, in the first panel we see the effect of the parameter $\tilde{p}$, finding that the increase of the marginal default probability shifts the distribution to the right. In the middle panel we notice that an increase of the value of $\omega$ amplifies the bimodality of the distribution separating the peaks. Finally, on the right we see that changes in the value of $\mu$ affect mostly the second peak of the distribution without affecting significantly the main peak. Overall, the contagion model allows to account for flexible shapes of the loss distribution. We point out that there is the potential for identifiability issues due to the similar effects of the parameters $\mu$ and $\omega$ on the loss curve. Thus, in the empirical analysis, $\mu_{i}$ was held constant to ensure model parsimony.
@@
Next, we compare the loss distributions under different models: pure contagion (CON), one factor Gaussian copula (OFG), conditional model (COND), and mixture model (MIX). To make the comparison fair, we set the parameters of each distribution in such a way that the expected losses (EL) and the pairwise default correlation between any two assets ($\rho(Z_{i},Z_{j})$) are the same for each model. LGD is set to 100%, and the parameters for each distribution are reported in Table 1. In particular, we set a baseline parametrization for the CON model, and we adjusted the correlation parameter $\rho$ in the models. Default correlation is computed for CON, COND, and MIX models as described in Section 4.3, while for the OFG model we computed them analytically as described e.g. in Vasicek (2002).
@@
Figure 2: Loss distribution for different models with the parameters in Table 1. The horizontal axis represents the number of defaults and, although the maximum possible number is 125, only the range up to 40 is displayed for clarity.
@@ 7.1 Dataset
We consider synthetic CDO tranches of the iTraxx Europe index, that is the main reference for the European market and it is the largest in terms of CDS index tranches [^2] The notional traded in iTraxx Europe index tranches in 2024 was USD 99.1 billion, ranking above the CDX North America Investment Grade, despite the decrease in volumes compared to the previous year (see S&P Dow Jones Indices - 2024 & H1 2025 Fixed Income Index Products Report, https://www.spglobal.com/spdji/en/documents/commentary/market-commentary-fixed-income-index-products-annual-report-2024-2025-h1.pdf). Tranches are quoted in terms of upfront paid by the protection buyer to enter in a contract with 100 basis point coupon. The main index is instead listed in terms of par spread (i.e. the level of the premium that balances the expected value of the two legs of the contract). The quotes for contracts with 5 years maturity have been downloaded from LSEG Workspace for the index and for four tranches: Equity (0-3%), Junior Mezzanine (3-6%), Senior Mezzanine (6-12%), Senior (12-100%). For each day we also download from LSEG Workspace the local 5Y CDS spread for the 125 constituents of the index to compute the marginal default probabilities (used to estimate the marginal default probabilities $\tilde{p}_{i}$), the industry sector to define $\mu_{i}$ (for CON-BNK we consider companies whose industry sector is “Banking”, while for CON-FIN we consider all the companies belonging to the “Banking”, “Finance”, or “Insurance” industry sectors), and the bootstrapped EURIRS swap rates as the risk free spot rates. We assume a 40% recovery rate following market convention. Every 6 months a new series is released, so that the index composition reflects the market characteristics. We download the dataset with daily frequency from October 31, 2019 to May 30, 2025, and for each date we consider the most recent iTraxx series available. Overall we considered 1457 trading days, excluding 2.8% of the dates from the analysis due to missing quotes for the CDO tranches.
@@ 7.2 Pricing with contagion and mix models
For the calibration of the model we use MATLAB 2024a, minimizing an objective function consisting in the sum of root squared percentage error of the quotes (upfronts for the tranches, par-spread for the whole index). Since some of the quotes were very close to zero, we added a translation of 0.1 in the quotes used in the objective function to avoid denominator close or equal to zero to improve the numerical stability of the optimization. The objective function is minimized using fmincon from the Optimization Toolbox, and the initial parameters supplied to the solver are $\omega_{0}=\rho_{0}=\pi_{0}=0.5$. Each optimization of the COND model (the most computationally intensive, since the contagion loss distribution must be computed for each value of the conditioning variable) takes typically less than two minutes (single core) on a Macbook air with M1 processor and 16 GB of RAM. We bound all the parameters between 0.05 and 0.95 to improve the numerical stability, and, with the exception of the OFG model, in the majority of times we do not obtain corner solutions. [^3] The code used to estimate the loss curve and to price CDOs under the CON, OFG, COND, and MIX models is available from the corresponding author upon request, together with a toy dataset containing the required inputs in the appropriate format.
@@
Table 3: iTraxx tranche calibrated quotes. The quotes for the CDO tranches are upfronts (running spread equal to 100 bps) while the ones for the index are par spreads. The lowest Mean Absolute Error (MAE) is highlighted in bold.
@@
To study the models’ adaptive boundaries in different market environments, we report the average error in a set of crisis scenarios. In particular, we consider the following 6 scenarios: COVID-19 Sharp Shock (Mar 1 – Jun 1, 2020), COVID-19 Recovery & Adaptation (Jun 1 – Dec 31, 2020), Post-Pandemic (Jan 1 – Dec 31, 2021), Ukraine / Energy Crisis (Feb 24 – Aug 31, 2022), European Interest Rate Tightening (Jul 27 – Oct 27, 2022), Post-Crisis Normalization (Jan 1, 2023 – May 31, 2025), and we fit the model each day. Table 4 reports the average MAE of the days in each period and the corresponding volatility. We see that the MIX models are the best performing in each scenario. As we previously noticed from Figure 3, the OFG tends to perform relatively better in quiet periods, while the contagion models (CON) show lower errors in the harshest periods. We point out that such observation may have implications on the interpretation of the model: the market perception of the loss curve may be closer to the CON model in distressed period with high perceived probability of default clustering, while it may be closer to the one implied by the Gaussian model in quiter periods. Thus, the proposed parametrization of the models may be used to gauge the market perception of the stability of the market: a higher-than-usual level of $\pi$ in the MIX model, may for instance be studied as a market-based indicator of an increased fear of financial contagion. Still, the focus of this paper is on static loss-distribution comparison. Evaluating model performance in a genuine time-series forecasting setting (e.g., temporal clustering or regime-switching predictability) and formally study the application of the framework to systemic risk analysis is outside the scope of this work and is left for future research.
@@ 8 Conclusions
In this paper we presented a new contagion model for the estimation of the loss distribution of a credit portfolio. The model that we introduce is related to the seminal work of Davis and Lo (2001) and it can achieve reasonable performance with heterogeneous portfolios.
We then introduced two alternative models for the construction of the loss distribution that incorporate the effect of a common factor, either using a mixture distribution or by conditioning on a common Gaussian factor. We apply the model to the problem of pricing synthetic CDO tranches of the iTraxx index, calibrating several specifications of the model on multiple dates with satisfactory performance on the studied dataset. In particular the best performance are obtained using the three parameters mixture models. The models are computationally efficient, and the optimal parameters have a clear economic interpretation.
In addition to fixed-income trading desks, the model has potential applications in the monitoring of systemic risk, as it allows to extract market-based forward looking information on the possible manifestation of systemic events, distinguishing between contagion events and joint defaults due to common factors. We leave the exploration of policy applications and the testing of the empirical properties of the model in other markets (e.g. the CDX North America Investment Grade index) to future works.